OpenStudy (ttop0816):

Evaluate |-6| + 3|4| + |-5| - 0 · 6. 0 18 23 25

OpenStudy (comm.dan):

What do u think ur answer is @ttop0816?

OpenStudy (ttop0816):

23?

OpenStudy (mathstudent55):

First take each absolute value. Then use the order of operations.

OpenStudy (comm.dan):

Then, what is ur answer? @ttop0816

OpenStudy (mathstudent55):

Correct.

OpenStudy (comm.dan):

Correct @ttop0816

OpenStudy (ttop0816):

thank you! mind sorting out something else with me too?

OpenStudy (ttop0816):

Evaluate -8 · |2|. -10 10 -16 16 isnt it 16?

OpenStudy (comm.dan):

Yes, go ahead. I don't mind. I will be here to guide u. Yes

OpenStudy (comm.dan):

Sorry, my bad That is incorrect. What do u think the answer is?

OpenStudy (mathstudent55):

No. The -8 remains as -8. it has not absolute value sign. -8 * 2 = -16

OpenStudy (comm.dan):

That is what I meant to say @mathstudent55

OpenStudy (comm.dan):

Anymore questions? @ttop0816

OpenStudy (ttop0816):

ohhh

OpenStudy (ttop0816):

then for this question Evaluate |5 - 15|. it would be |13| right?

OpenStudy (ttop0816):

because they had the negative inside of the absolute value?

OpenStudy (comm.dan):

No.

OpenStudy (ttop0816):

oh sorry i meant 10!

OpenStudy (comm.dan):

Yes

OpenStudy (comm.dan):

I think what is 15-5 and then what is the opposite of that number

OpenStudy (mathstudent55):

|5 - 15| = |-10| = 10

OpenStudy (comm.dan):

That is what I do, too. @mathstudent55

OpenStudy (ttop0816):

then what about this one? Evaluate 3|-3|. would that still be -9?

OpenStudy (comm.dan):

No

OpenStudy (comm.dan):

Think what is the opposite of -3 and think of that every time u have a negative number in absolute value terms

OpenStudy (ttop0816):

so it would be just 9?

OpenStudy (comm.dan):

 @ttop0816

OpenStudy (comm.dan):

Yes

OpenStudy (comm.dan):



OpenStudy (mathstudent55):

First you need to take the absolute value of -3, then you multiply it by 3: 3|-3| = 3 * 3 = 9